Propose a possible structure for the compounds that produce the following spectrums. The molecular formula is given. Justify your answer identifying the functional groups present.

CH10 761 2909 946 2940 2964 136€ 1173 ON H5 1161 CH.02 17963 976 1466 3418 1157 1379 2967 1724
Answer #1

OCH HE CORE = 4 + 2 = 5-4 = 1! IR = 2964 cmt sp3 C-H stretch 1718 cm carbonyl stretch & cz Hy No (OBE = 3 - (2-1) + 1 = 4-3=